# Uniformly-$S$-pseudo-projective modules

### v1

@@ Abstract.
In this paper, we introduce the notion of uniformly-$S$-pseudo-projective ($u$-$S$-pseudo-projective) modules as a generalization of $u$-$S$-projective modules. Let $R$ be a ring and $S$ a multiplicative subset of $R$. An $R$-module $P$ is said to be $u$-$S$-pseudo-projective if for any submodule $K$ of $P$, there is $s\in S$ such that for any $u$-$S$-epimorphism $f:P\to\frac{P}{K}$, $sf$ can be lifted to an endomorphism $g:P\to P$. Some properties of this notion are obtained. For example, we prove that an $R$-module $M$ is $u$-$S$-quasi-projective if and only if $M\oplus M$ is $u$-$S$-pseudo-projective. A new characterization of $u$-$S$-semisimple rings is given in terms of this notion.
@@ 1. Introduction
In this paper, all rings are commutative with a nonzero identity, and all modules are unitary. Let $R$ be a ring. A subset $S$ of $R$ is called a multiplicative subset of $R$ if $1\in S$, $0\notin S$, and $s_{1}s_{2}\in S$ for all $s_{1},s_{2}\in S$. Let $S$ be a multiplicative subset of a ring $R$ and $M,N,L$ be $R$-modules.
@@
In Section 2, we study some properties of $u$-$S$-pseudo-projective modules. Firstly, we show in Remark 2.2 (1) that the notions of $u$-$S$-pseudo-projective modules and pseudo-projective modules coincide under the condition that every element of $S$ is a unit of $R$. Corollary 2.8 shows that an $R$-module $M$ is $u$-$S$-quasi-projective if and only if $M\oplus M$ is $u$-$S$-pseudo-projective. Theorem 2.15 gives a new characterization of $u$-$S$-semisimple rings in terms of $u$-$S$-pseudo-projective modules. Theorem 2.17 characterizes rings in which every $u$-$S$-pseudo-projective module is $u$-$S$-projective.
@@ 2. $u$-$S$-pseudo-projective modules
Let $S$ be a multiplicative subset of a ring $R$.
@@
($\Leftarrow$). For $n=2$, apply Corollary 2.8. For $n>2$, since $M^{(2)}\oplus M^{(n-2)}\cong M^{(n)}$ is $u$-$S$-pseudo-projective, then by Proposition 2.3 (2), $M^{(2)}$ is $u$-$S$-pseudo-projective. Thus $M$ is $u$-$S$-quasi-projective by Corollary 2.8.
Corollary 2.10.
Proof.
@@
Let $R=\mathbb{Z}$ and $S=\{1,2,3,\cdots\}$. Let $\mathbb{P}$ be the set of all prime numbers and $M:=\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}$. Then $M$ is a $u$-$S$-quasi-projective module that is not $u$-$S$-projective by [2, Example 3.3]. By Corollary 2.8, $M\oplus M$ is a $u$-$S$-pseudo-projective module. However, $M\oplus M$ is not $u$-$S$-projective by [2, Corollary 2.8 (1)].
@@
$(1)\Rightarrow(2)$. Follows from Corollary 2.10.
@@
The following theorem gives a new characterization of $u$-$S$-semisimple rings in terms of $u$-$S$-pseudo-projective modules.
Theorem 2.15.
Let $S$ be a multiplicative subset of a ring $R$. Then $R$ is $u$-$S$-semisimple if and only if every $R$-module is $u$-$S$-pseudo-projective.
Proof.
Let $R$ be a $u$-$S$-semisimple ring. Then every $R$-module is $u$-$S$-quasi-projective by [2, Theorem 3.11]. So by Remark 2.2 (2), every $R$-module is $u$-$S$-pseudo-projective. Conversely, let $M$ be any $R$-module. By hypothesis, $M\oplus M$ is $u$-$S$-pseudo-projective. So by Corollary 2.8, $M$ is $u$-$S$-quasi-projective. Hence every $R$-module is $u$-$S$-quasi-projective. Therefore, $R$ is $u$-$S$-semisimple by [2, Theorem 3.11]. ∎
Theorem 2.16.
  1. (1)
    Every $u$-$S$-quasi-projective module is $u$-$S$-projective.
  2. (2)
Proof.
$(1)\Rightarrow(2)$. Let $M$ and $N$ be two $u$-$S$-quasi-projective modules. Then by (1), $M$ and $N$ are $u$-$S$-projective modules. So $M\oplus N$ is $u$-$S$-projective by [9, Proposition 2.14 (1)]. Hence $M\oplus N$ is $u$-$S$-quasi-projective by Remark 2.2 (2).
Theorem 2.17.
  1. (1)
    Every $u$-$S$-pseudo-projective module is $u$-$S$-projective.
  2. (2)
Proof.
$(1)\Rightarrow(2)$. Let $M$ and $N$ be two $u$-$S$-pseudo-projective modules. Then $M$ and $N$ are $u$-$S$-projective modules and so $M\oplus N$ is $u$-$S$-projective. Hence $M\oplus N$ is $u$-$S$-pseudo-projective by Remark 2.2 (2).
$(2)\Rightarrow(1)$. We have two claims:
Claim 1: every $u$-$S$-pseudo-projective is $u$-$S$-quasi-projective. Let $M$ be a $u$-$S$-pseudo-projective module. Then $M\oplus M$ is $u$-$S$-pseudo-projective by (2). So by Corollary 2.8, $M$ is $u$-$S$-quasi-projective.
Claim 2: every $u$-$S$-quasi-projective is $u$-$S$-projective. Let $M$ and $N$ be two $u$-$S$-quasi-projective modules. Then $M$ and $N$ are $u$-$S$-pseudo-projective modules by Remark 2.2 (2). By (2), $M\oplus N$ is $u$-$S$-pseudo-projective and hence by Claim 1, $M\oplus N$ is $u$-$S$-quasi-projective. So every direct sum of two $u$-$S$-quasi-projective modules is $u$-$S$-quasi-projective. Thus, by Theorem 2.16, every $u$-$S$-quasi-projective is $u$-$S$-projective. Therefore, by Claims 1 and 2, every $u$-$S$-pseudo-projective module is $u$-$S$-projective.
Acknowledgment: The authors would like to thank the referees for their valuable comments.

### v2

@@ Abstract.
In this paper, we introduce the notion of uniformly-$S$-pseudo-projective ($u$-$S$-pseudo-projective) modules as a generalization of $u$-$S$-projective modules. Let $R$ be a ring and $S$ a multiplicative subset of $R$. An $R$-module $P$ is said to be $u$-$S$-pseudo-projective if for any submodule $K$ of $P$, there is $s\in S$ such that for any $u$-$S$-epimorphism $f:P\to\frac{P}{K}$, $sf$ can be lifted to an endomorphism $g:P\to P$. Some properties of this notion are obtained. For example, we prove that an $R$-module $M$ is $u$-$S$-quasi-projective if and only if $M\oplus M$ is $u$-$S$-pseudo-projective. Also, we prove that if $A\oplus B$ is $u$-$S$-pseudo-projective, then any $u$-$S$-epimorphism $f:A\to B$ $u$-$S$-splits. Several new characterizations of $u$-$S$-semisimple rings are given.
@@ 1. Introduction
In this paper, all rings are commutative with a nonzero identity, and all modules are unitary unless otherwise stated. Let $R$ be a ring. A subset $S$ of $R$ is called a multiplicative subset of $R$ if $1\in S$, $0\notin S$, and $s_{1}s_{2}\in S$ for all $s_{1},s_{2}\in S$. Let $S$ be a multiplicative subset of a ring $R$ and $M,N,L$ be $R$-modules.
@@
In Section 2, we study some properties of $u$-$S$-pseudo-projective modules. Firstly, we show in Remark 2.2 (1) that the notions of $u$-$S$-pseudo-projective modules and pseudo-projective modules coincide under the condition that every element of $S$ is a unit of $R$. In Corollary 2.10, we show that an $R$-module $M$ is $u$-$S$-quasi-projective if and only if $M\oplus M$ is $u$-$S$-pseudo-projective. In Proposition 2.13, we show that if $M$ is an $R$-module and $f:P\to M$ is an epimorphism with $P$ projective, then $M$ is $u$-$S$-projective if and only if $P\oplus M$ is $u$-$S$-pseudo-projective. Theorem 2.15 gives several new characterizations of $u$-$S$-semisimple rings. Theorem 2.17 characterizes rings in which every $u$-$S$-pseudo-projective module is $u$-$S$-projective.
@@ 2. $u$-$S$-pseudo-projective modules
Let $S$ be a multiplicative subset of a ring $R$. Then the following statements hold:
@@
Let $S$ be a multiplicative subset of a ring $R$. Recall that $R$ is called $S$-von Neumann regular if for any $a\in R$, there exist $s\in S$ and $r\in R$ such that $sa=ra^{2}$ [10]. Equivalently, if the ring $R_{S}$ is von Neumann regular [10, Proposition 3.10]. Next, we extend the notion of $S$-von Neumann regular rings to noncommutative rings.
Definition 2.6.
Let $R$ be a noncommutative ring and $S$ a multiplicative subset of $R$. $R$ is said to be $S$-von Neumann regular if for any $a\in R$, there exist $s\in S$ and $r\in R$ such that $sa=ara$.
The following theorem gives sufficient conditions for the ring $\text{End}_{R}(P)$ to be $S$-von Neumann regular.
Theorem 2.7.
Let $S$ be a multiplicative subset of a ring $R$ and let $P$ be a $u$-$S$-pseudo-projective module such that $\text{Im}(e)$ is a direct summand of $P$ for every $e\in\text{End}_{R}(P)$. Then $\text{End}_{R}(P)$ is $S$-von Neumann regular ring.
Proof.
Let $e\in\text{End}_{R}(P)$. Then $\text{Im}(e)$ is a direct summand of $P$. By the proof of Corollary 2.5, the exact sequence $0\to\text{Ker}(e)\to P\xrightarrow{\eta}\frac{P}{\text{Ker}(e)}\to 0$ is $u$-$S$-split. So there is $s\in S$ and $R$-homomorphism $v:\frac{P}{\text{Ker}(e)}\to P$ such that $\eta v=s1_{\frac{P}{\text{Ker}(e)}}$. For $x\in P$, let $\overline{x}=x+\text{Ker}(e)$. Then $s\overline{x}=\eta v(\overline{x})=\overline{v(\overline{x})}$. So $sx-v(\overline{x})\in\text{Ker}(e)$, It follows that $sP\subseteq\text{Ker}(e)+\text{Im}(v)$. Also, we have $s\big(\text{Ker}(e)\cap\text{Im}(v)\big)=0$. Indeed, if $v(\overline{y})\in\text{Ker}(e)$, then $s\overline{y}=\eta v(\overline{y})=\overline{v(\overline{y})}=\overline{0}$ and so $sv(\overline{y})=0$. Let $P=A\oplus\text{Im}(e)$ be an internal direct sum. We have $se(P)\subseteq e(\text{Im}(v))$. Define $q:P\to P$ by $q(a+e(x))=sb$ where $a\in A$, $x\in P$, and $b\in\text{Im}(v)$ is such that $se(x)=e(b)$. Then $q$ is well defined. Indeed, if $a+e(x)=a^{\prime}+e(x^{\prime})$, then $a-a^{\prime}=e(x^{\prime}-x)\in A\cap\text{Im}(e)=0$. So $a=a^{\prime}$ and $x^{\prime}-x\in\text{Ker}(e)$. Write $se(x)=e(b)$ and $se(x^{\prime})=e(b^{\prime})$ for some $b,b^{\prime}\in\text{Im}(v)$. Since $e(b^{\prime}-b)=se(x^{\prime}-x)=0$, then $b^{\prime}-b\in\text{Ker}(e)\cap\text{Im}(v)$ and so $s(b^{\prime}-b)=0$. Hence $sb=sb^{\prime}$. Finally, we show that $s^{2}e=eqe$. Let $x\in P$. So $se(x)=e(b)$ for some $b\in\text{Im}(v)$. Thus $eqe(x)=e(q(0+e(x)))=e(sb)=se(b)=s^{2}e(x)$. Since $x\in P$ was arbitrary, we have $s^{2}e=eqe$. Therefore, $\text{End}_{R}(P)$ is $S$-von Neumann regular ring. ∎
@@
($\Leftarrow$). For $n=2$, apply Corollary 2.10. For $n>2$, since $M^{(2)}\oplus M^{(n-2)}\cong M^{(n)}$ is $u$-$S$-pseudo-projective, then by Proposition 2.3 (2), $M^{(2)}$ is $u$-$S$-pseudo-projective. Thus $M$ is $u$-$S$-quasi-projective by Corollary 2.10. ∎
@@
Let $R=\mathbb{Z}$ and $S=\{1,2,3,\cdots\}$. Let $\mathbb{P}$ be the set of all prime numbers. Then $\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}$ is a $u$-$S$-quasi-projective module that is not $u$-$S$-projective by [2, Example 3.3]. By Corollary 2.10, $\Big(\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}\Big)\oplus\Big(\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}\Big)$ is a $u$-$S$-pseudo-projective module. However, $\Big(\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}\Big)\oplus\Big(\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}\Big)$ is not $u$-$S$-projective by [2, Corollary 2.8 (1)].
Proposition 2.13.
Let $S$ be a multiplicative subset of a ring $R$. The following statements hold:
  1. (1)
    If $A\oplus B$ is $u$-$S$-pseudo-projective, then any $u$-$S$-epimorphism $f:A\to B$ $u$-$S$-splits.
  2. (2)
    If $M$ is an $R$-module and $f:P\to M$ is an epimorphism with $P$ projective, then $M$ is $u$-$S$-projective if and only if $P\oplus M$ is $u$-$S$-pseudo-projective.
Proof.
  1. (1)
    Let $f:A\to B$ be any $u$-$S$-epimorphism. Since $A\oplus B$ is $u$-$S$-pseudo-projective, then by Proposition 2.9, $B$ is $u$-$S$-projective relative to $A$ and hence by [2, Lemma 2.15], the $u$-$S$-exact sequence $0\to\text{Ker}(f)\to A\xrightarrow{f}B\to 0$ $u$-$S$-splits. Thus $f$ $u$-$S$-splits.
  2. (2)
    Suppose that $M$ is $u$-$S$-projective. Since $P$ is projective, then by [9, Corollary 2.11], $P$ is $u$-$S$-projective. So by [9, Proposition 2.14 (1)], $P\oplus M$ is $u$-$S$-projective and hence by Remark 2.2 (2), $P\oplus M$ is $u$-$S$-pseudo-projective. Conversely, suppose that $P\oplus M$ is $u$-$S$-pseudo-projective. Then by part (1) and since $f:P\to M$ is a $u$-$S$-epimorphism, we have $f$ $u$-$S$-splits. So the exact sequence $0\to\text{Ker}(f)\to P\xrightarrow{f}M\to 0$ $u$-$S$-splits. Hence by [5, Lemma 2.8], $P$ is $u$-$S$-isomorphic to $\text{Ker}(f)\oplus M$. Since $P$ is $u$-$S$-projective, then by [9, Proposition 2.14 (3)] and [2, Corollary 2.8 (1)], $M$ is $u$-$S$-projective.
For an $R$-module $M$, let $E(M)$ denote the injective envelope of $M$.
Proposition 2.14.
Let $S$ be a multiplicative subset of a ring $R$ and $M$ an $R$-module. Then $M$ is $u$-$S$-injective if and only if $M\oplus E(M)$ is $u$-$S$-pseudo-injective.
Proof.
Let $M$ be $u$-$S$-injective. Then $M\oplus E(M)$ is $u$-$S$-injective by [6, Proposition 4.7]. So by [3, Remark 2.4 (2)], $M\oplus E(M)$ is $u$-$S$-pseudo-injective. Conversely, suppose that $M\oplus E(M)$ is $u$-$S$-pseudo-injective. Then by [3, Theorem 2.6], the monomorphism $i_{M}:M\to E(M)$ $u$-$S$-splits. So the exact sequence $0\to M\to E(M)\to\frac{E(M)}{M}\to 0$ $u$-$S$-splits. Hence by [5, Lemma 2.8], $E(M)$ is $u$-$S$-isomorphic to $M\oplus\frac{E(M)}{M}$. But $E(M)$ is $u$-$S$-injective, so by [6, Proposition 4.7 (3)] and [2, Corollary 2.8 (2)], $M$ is $u$-$S$-injective. ∎
The following theorem gives several characterizations of $u$-$S$-semisimple rings.
Theorem 2.15.
  1. (1)
    $R$ is $u$-$S$-semisimple.
  2. (2)
    Every $R$-module is $u$-$S$-pseudo-projective.
  3. (3)
    Every direct sum of an $R$-module and a free $R$-module is $u$-$S$-pseudo-projective ($u$-$S$-pseudo-injective).
  4. (4)
    Every direct sum of an $R$-module and a free $R$-module is $u$-$S$-quasi-projective ($u$-$S$-quasi-injective).
Proof.
$(1)\Rightarrow(2)$. Let $R$ be a $u$-$S$-semisimple ring. Then every $R$-module is $u$-$S$-quasi-projective by [2, Theorem 3.11]. So by Remark 2.2 (2), every $R$-module is $u$-$S$-pseudo-projective.
$(2)\Rightarrow(1)$. Let $M$ be any $R$-module and let $f:P\to M$ be an epimorphism with $P$ projective. Then by (2), $P\oplus M$ is $u$-$S$-pseudo-projective. So by Proposition 2.13 (2), $M$ is $u$-$S$-projective. Hence every $R$-module is $u$-$S$-projective. Therefore, $R$ is $u$-$S$-semisimple by [9, Theorem 3.5].
$(1)\Rightarrow(3)$. This follows from $(1)\Rightarrow(2)$ and [3, Theorem 2.14].
$(3)\Rightarrow(1)$. Let $F$ be any free $R$-module. Then by (3), for any $R$-module $M$, $M\oplus F$ is $u$-$S$-pseudo-projective ($u$-$S$-pseudo-injective). By Proposition 2.9 ([3, Proposition 2.13]), $M$ is $u$-$S$-projective ($u$-$S$-injective) relative to $F$ for any $R$-module $M$. Thus every $R$-module is $u$-$S$-projective ($u$-$S$-injective) relative to $F$, By [2, Theorem 2.16], $F$ is $u$-$S$-semisimple. Hence every free $R$-module is $u$-$S$-semisimple. Therefore, $R$ is $u$-$S$-semisimple.
$(1)\Leftrightarrow(4)$. This follows from [2, Theorem 2.16, Proposition 3.8, and Theorem 3.11]. ∎
Theorem 2.16.
  1. (1)
    Every $u$-$S$-quasi-projective module is $u$-$S$-projective.
  2. (2)
    Every direct sum of two $u$-$S$-quasi-projective modules is $u$-$S$-projective.
  3. (3)
Proof.
$(1)\Rightarrow(2)$. Let $M$ and $N$ be two $u$-$S$-quasi-projective modules. Then by (1), $M$ and $N$ are $u$-$S$-projective modules. So $M\oplus N$ is $u$-$S$-projective by [9, Proposition 2.14 (1)].
$(2)\Rightarrow(3)$. This is clear from Remark 2.2 (2).
Theorem 2.17.
  1. (1)
    Every $u$-$S$-pseudo-projective module is $u$-$S$-projective.
  2. (2)
    Every direct sum of two $u$-$S$-pseudo-projective modules is $u$-$S$-projective.
  3. (3)
Proof.
$(1)\Rightarrow(2)$. Let $M$ and $N$ be two $u$-$S$-pseudo-projective modules. Then $M$ and $N$ are $u$-$S$-projective modules and so $M\oplus N$ is $u$-$S$-projective.
$(2)\Rightarrow(3)$. This follows from Remark 2.2 (2).
$(3)\Rightarrow(1)$. Let $M$ be any $u$-$S$-pseudo-projective module. Let $f:P\to M$ be an epimorphism with $P$ projective. Then by (3), $P\oplus M$ is $u$-$S$-pseudo-projective. Thus $M$ is $u$-$S$-projective by Proposition 2.13 (2).
@@
$(1)\Rightarrow(2)$. Follows from Remark 2.19.
@@
Acknowledgments: The authors would like to thank the referees for their valuable comments.